KS123:

does anyone know this x3+y3+z3=k?

BBYGIRL:

what the problem look like

Astro:

@KS123 Do you have a screenshot of the problem?

BBYGIRL:

yuh

OronSH:

There exists an integer solution to \(x^3+y^3+z^3=k\) for all \(|k|\leq100\).

jhonyy9:

1 +8 + 27 = k do you think on this ? or 8 +27 +64 = k or ... try get again you

Crasher:

x+x+x+y+y+y+z+z+z+t+t=k

bubbly1belle:

what letter are you solving for?

surjithayer:

\[\left| k \right|\le 100 \implies -100\le k \le 100\] 1. 0+0+0=0 0+0+1=1 0+1+0=1 1+0+0=1 0+1+1=2 1+0+1=2 1+1+0=2 2. 1+8+27=36 0+8+27=35 -1+8+27=34 -8+8+27=27 -27+8+27=8 -64+8+27=-29 -125+8+27=-90 3. .... you can select many solutions.